OpenStudy (aziza3):

The horizontal component of a velocity vector is 23 m/s. If the object travels at an angle of 87° to the horizontal, how fast is it moving?

imqwerty (imqwerty):

\(Horizontal~component=v cos\theta\) \(23=v \times cos87° \) :) just put the value of \(cos87\) and solve to get the value of \(v\) which is ur answer

OpenStudy (shamim):

U know Horizontal component=v*cos theta Vertical component=v*sin theta